In a film, a character is criticized for marrying a woman when he is three times her age. He witty replies "ah but in 24 years time I will only be twice her age". How old are they
Answer by Cromlix(4381)

Hi there,
he is three times her age.
M = 3W
ah but in 24 years time I will only be twice her age".
M + 24 = 2(W + 24)
Replace M with 3W from top equation.
3W + 24 = 2W + 48
Collect like terms.
3W - 2W = 48 - 24
W = 24
Woman is 24 years old.
Man is 72 years old.
Hope this helps :-)